OpenStudy (anonymous):

Find the slope of the tangent line to the curve 2x^4-3x^2+sq.x-5/x^2 at point (1,-5)

OpenStudy (amistre64):

is the x^2 the denominator of everything else?

OpenStudy (anonymous):

no just the denominator of 5

OpenStudy (amistre64):

ok...so the last term is 5/x^2 everything else is as is..

OpenStudy (anonymous):

i was thinking about rewriting it as 5x^-2, but i'm not sure

OpenStudy (anonymous):

correct

OpenStudy (amistre64):

go ahead, its all good :)

OpenStudy (anonymous):

ook cool

OpenStudy (amistre64):

8x^3 -6x +(1/2sqrt(x)) +(10/x^3) is what i get... maybe its right?

OpenStudy (anonymous):

yea i got 8x^3-6x+1/2x^-1/2-10x^-3, should i substitute after this?

OpenStudy (amistre64):

yep, but I think you only need the "x" value

OpenStudy (amistre64):

8 -6 + (1/2) +10

OpenStudy (anonymous):

yea so just substitute the x value

OpenStudy (amistre64):

and youve got a typo,,,that last term should be a (+); -5(-2)=10

OpenStudy (anonymous):

oh yea thanks

OpenStudy (amistre64):

I get 6 as the slope... if I did it right :)

OpenStudy (anonymous):

uhm i'm not sure if we're doing it right cuz the problem is actually asking me for the equation of the line

OpenStudy (amistre64):

the derivative is the equation of the slope of the line; or maybe it should be in y = mx+b format?

OpenStudy (anonymous):

yea i think that's what they mean y=mx+b format

OpenStudy (amistre64):

y = 6x +b at (x=1,y=-5)?? -5 = 6(1) + b -5 = 6 + (-11) y = 6x-11 then

OpenStudy (anonymous):

hold up, how did u get 6 as the slope, i'm up to 8-6+1/2+10

OpenStudy (amistre64):

f'(x) = m 8-6=2+10=12...ahhh, I see; I multiplied by 1/2

OpenStudy (amistre64):

12.5 happy :)

OpenStudy (anonymous):

ook cool

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

-5 = 12.5(1) + b -5 -12.5 = b -17.5 = b y = 12.5 - 17.5...

OpenStudy (amistre64):

squeeze an "x" right up next to that 12.5

OpenStudy (amistre64):

or do they want it in Ax + By = C format?

OpenStudy (anonymous):

y=mx+b format

OpenStudy (amistre64):

you can use the line equation to figure out a new "x value" when iterating a function solve it...

OpenStudy (anonymous):

uhm not sure about that... does that mean that y=6x-11 is not it

OpenStudy (amistre64):

yeah, since i miscalculated the first time with a "6", its a bad apple...toss it out before it spoils the rest :)

OpenStudy (anonymous):

ook

OpenStudy (anonymous):

ok then it's 8-6+1/2+10=12+1/2=25/2

OpenStudy (anonymous):

so the slope is 25/2

OpenStudy (anonymous):

got it thanks :)